OpenStudy (anonymous):

Solve for x 4x^2=-4x+3

OpenStudy (er.mohd.amir):

4x2+4x-3=0 4x2+(6-2)x-3=0 4x2+6x-2x-3=0

OpenStudy (er.mohd.amir):

2x(x+3)-1(2x+3)=0 (2x-1)(2x+3)=0 x=1/2 or -3/2